# Reinforcement Learning on AYA Dyads to Enhance Medication Adherence

### v1

@@ Abstract
Medication adherence is critical for the recovery of adolescents and young adults (AYAs) who have undergone hematopoietic cell transplantation (HCT). However, maintaining adherence is challenging for AYAs after hospital discharge, who experience both individual (e.g. physical and emotional symptoms) and interpersonal barriers (e.g., relational difficulties with their care partner, who is often involved in medication management). To optimize the effectiveness of a three-component digital intervention targeting both members of the dyad as well as their relationship, we propose a novel Multi-Agent Reinforcement Learning (MARL) approach to personalize the delivery of interventions. By incorporating the domain knowledge, the MARL framework, where each agent is responsible for the delivery of one intervention component, allows for faster learning compared with a flattened agent. Evaluation using a dyadic simulator environment, based on real clinical data, shows a significant improvement in medication adherence (approximately 3%) compared to purely random intervention delivery. The effectiveness of this approach will be further evaluated in an upcoming trial.
@@ 1 Introduction
The use of MARL decouples the decision processes of different intervention components, thus improving interpretability of the agent model design. This improved interpretability allows us to incorporate domain knowledge into the agent-specific algorithm designs to address challenge (2). To further accelerate learning, we propose a novel reward engineering method that learns a less noisy surrogate reward function for each component. Through evaluation in a carefully designed dyadic environment, we demonstrate both the superior performance of our proposed algorithm and strong collaborative behavior among the three agents. Lack of collaboration is often a critical issue in MARL [16].
@@ 2 RL Framework and Domain Knowledge
HCT treatment is followed by an outpatient 14-weeks twice-daily medication regimen. Decision times within the 14 weeks are denoted by $(w,d,t)$ where $w\in\{1,\dots,14\}$ is the week index, $d\in\{1,\dots,7\}$ is the day index, and $t\in\{1,2\}$ is the decision window within a day.
@@
Action space. All actions are binary (deliver versus do not deliver intervention content); see Tables 1,2. When the current time $(d=1,t=1)$ is the first decision time on the first day of the week, the agent chooses a three-dimensional action corresponding to all three interventions components. If the current time is the first time on a day after the first day of the week $(d>1,t=1)$, the agent chooses a two-dimensional action corresponding to only the AYA intervention and the care partner components. At the second time on each day ($t=2$) the agent chooses a one-dimensional action corresponding to only the AYA intervention component.
Observation space. Apart from the dynamic action space, we collect observations about different components at different time scales as well; see Table 2. At each time $(w,d,t)$, we collect the current medication adherence and digital intervention burden from the AYA component. In the end of each day $d$, we collect the psychological distress and digital intervention burden from the care partner component. In the end of a week $w$, we collect the relationship quality from questionnaires from both the AYA and the care partner.
@@ 3 Domain Knowledge through Causal Diagram
Our algorithm design is guided by domain knowledge encoded as the causal diagram in Fig. 1. This diagram describes the scientific team’s understanding of the primary causal relationships between the variables in each component listed in Table 2. Note that the causal relationships are likely more complex and direct paths may exist between any two variables. However the scientific team believes that these other paths are likely to be less detectable given the noise in digital intervention data. We summarize the primary pathways that interventions can take to effect the AYA’s adherence in the following.
@@
    Game intervention. The game intervention $A_{w}^{\tiny\texttt{REL}}$ has two pathways by which it is expected to effect AYA’s adherence. First, $A_{w}^{\tiny\texttt{REL}}$ is expected to increase the AYA’s burden $B_{w,d,t}^{\tiny\texttt{AYA}}$ throughout the week $w$. And AYA’s burden $B_{w,d,t}^{\tiny\texttt{AYA}}$ is expected to decrease the AYA’s adherence $R_{w,d,t}^{\tiny\texttt{AYA}}$ (blue arrows). Second, the game intervention $A_{w}^{\tiny\texttt{REL}}$ is expected to effect next week AYA’s adherence $R_{w+1,d,t}^{\tiny\texttt{AYA}}$ by improving the end of the week relationship quality $Y_{w}^{\tiny\texttt{REL}}$ (green arrows).
@@
    Care partner intervention. The care-partner intervention $A_{w,d}^{\tiny\texttt{CARE}}$ is expected to effect the AYA’s adherence indirectly. First, $A_{w,d}^{\tiny\texttt{CARE}}$ should decrease the care partner’s psychological distress $Y_{w,d}^{\tiny\texttt{CARE}}$, which should increase the end of week relationship quality $Y_{w}^{\tiny\texttt{REL}}$ (yellow arrows). Second, $A_{w,d}^{\tiny\texttt{CARE}}$ should increase the care partner’s burden $B_{w,d}^{\tiny\texttt{CARE}}$, which should decrease the the end of week relationship quality $Y_{w}^{\tiny\texttt{REL}}$ (blue arrows).
We further note that the variables from different components are generally independent conditioned on the bottleneck variables, e.g., the relationship quality that blocks all the paths from the care partner variables to the AYA’s adherence. This forms the basis of our multi-agent RL design.
@@ 4 Proposed Multi-Agent RL Approach
The conditional independence property observed from Fig. 1 motivates us to design a multi-agent RL (MARL) comprising three agents: the AYA agent, the care partner agent, and the relationship agent. Each makes decisions at different time scales for their own component.
@@ 4.1 Surrogate Reward Function Design Through Domain Knowledge
Typical MARL [16] with independent learners considers agents making decisions without communication. In our study, the lack of communication is due to the different time scales–the relationship agent that makes decisions in the beginning of a week may not predict the AYA and care partner agents’ decisions throughout the week. This may prevent the agents from collaborating. For example, the relationship agent may choose to always intervene so as to improve the relationship quality (the primary goal of the game intervention), which may not be optimal for the AYA’s adherence.
@@
Furthermore, the effects of care partner intervention and the game intervention are highly delayed. The game intervention improves end of week relationship quality with a significant delayed effect onto the adherence in the next week. The care partner intervention (positive messages for the care partner) is designed to mitigate the care partner’s psychological distress, which only has indirect and delayed effects on the AYA’s adherence.
To address the above two issues, we engineer the reward function to account for the delayed effects and across-component effects of each intervention component to promot collaboration. Similar reward engineering in the context of digital interventions is discussed in [30]. Our approach is distinct in that we explore the principles for incorporating domain knowledge to guide the reward function design.
@@
Domain knowledge informed surrogate reward functions. We introduce the surrogate reward functions for the relationship agent and the care partner agent. As informed by Fig. 1, the delayed effect of the game intervention is through the relationship quality and the AYA burden. This motivates us to fit a linear model to predict the sum of medication adherence within week $w$, $\sum_{d=1}^{7}\sum_{t=1}^{2}R_{w,d,t}^{\tiny\texttt{AYA}}$, using $(1,Y_{w-1}^{\tiny\texttt{REL}},{B}_{w,1,1}^{\tiny\texttt{AYA}},A_{w}^{\tiny\texttt{REL}},A_{w}^{\tiny\texttt{REL}}\cdot Y_{w-1}^{\tiny\texttt{REL}})$ as the covariates. To account for the delayed effect, we engineer the surrogate reward function for the relationship agent as: $r_{w}^{\tiny\texttt{REL}}=(1,Y_{w-1}^{\tiny\texttt{REL}},{B}_{w,1,1}^{\tiny\texttt{AYA}},A_{w}^{\tiny\texttt{REL}},A_{w}^{\tiny\texttt{REL}}\cdot Y_{w-1}^{\tiny\texttt{REL}}){\bm{\beta}}^{\tiny\texttt{REL}}+\max_{a}(1,Y_{w}^{\tiny\texttt{REL}},{B}_{w+1,1,1}^{\tiny\texttt{AYA}},a,a\cdot Y_{w}^{\tiny\texttt{REL}}){\bm{\beta}}^{\tiny\texttt{REL}},$ where ${\bm{\beta}}^{\tiny\texttt{REL}}\in\mathbb{R}^{5}$ are Bayesian linear regression estimates.  [^1] We choose the prior mean to reflect our guesses on the sign the coefficients. The prior variance is chosen so the prior mean dominates until around the 5th dyads. The complete prior is provided in Appendix. The above reward yields a two-step greedy policy, which is a good enough approximation for the total sum of the medication adherence. We opt for a simple, linear model here because the bias trade-off is justified by the faster learning and reduction in noise.
@@
The design of the care partner agent is similar. A key observation is that the end of the week relationship quality blocks all the paths from the care partner variables to the AYA’s adherence. Thus, we fit a linear model to predict the end of week relationship quality $Y_{w+1}^{\tiny\texttt{REL}}$ using $(1,Y_{w,d}^{\tiny\texttt{CARE}},B_{w,d+1}^{\tiny\texttt{CARE}},Y_{w-1}^{\tiny\texttt{REL}},A_{w,d}^{\tiny\texttt{CARE}})$ as covariates. The surrogate reward function is: $r_{w,d}^{\tiny\texttt{CARE}}=(1,Y_{w,d}^{\tiny\texttt{CARE}},B_{w,d+1}^{\tiny\texttt{CARE}},Y_{w-1}^{\tiny\texttt{REL}},A_{w,d}^{\tiny\texttt{CARE}}){\bm{\beta}}^{\tiny\texttt{CARE}}$, where ${\bm{\beta}}^{\tiny\texttt{CARE}}\in\mathbb{R}^{5}$ are Bayesian linear regression estimates.
@@ 5 Results
Our environment is based on Roadmap 2.0 dataset involving 171 dyads, each consisting of a patient undergone HCT (target person) and a care partner. Roadmap 2.0 provides daily positive psychology interventions to the care partner only. Roadmap 2.0 collects wearable devices data, for example, physical activity, and self-report data, for example, mood score.
We build upon the environment design in [11], which also uses the Roadmap 2.0 data, but primarily focuses on AYA and relationship intervention component. We extend the environment to include the care partner intervention component. Specifically, we fit a separate multi-variate linear model for each component’s outcome (ie., $R_{w,d,t}^{\tiny\texttt{AYA}},Y_{w,d}^{\tiny\texttt{CARE}},Y_{w}^{\tiny\texttt{REL}}$) in the dataset. These models simulate the user trajectories under no intervention.
To simulate outcomes under treatments, we impute the treatment effects of the interventions and the effects of app burden, so the induced standard treatment effects (STE)  [^2] STE here is defined as the difference in the mean of the primary outcomes under the proposed intervention package and these under no intervention, which is further standardized by the standard deviation under no intervention. are around 0.15, 0.3, and 0.5. These STEs are commonly seen in behavioral science studies [3]. A complete description and code of the dyadic environment is provided in supplementary material  [^3] https://github.com/StatisticalReinforcementLearningLab/ADAPTS-HCT-AIME.
@@ 5.1 Cumulative Adherence Improvement
We simulate 25 dyads, the planned sample size in the upcoming pilot study, by sampling dyads sequentially with replacement from Roadmap 2.0 dataset. Each dyad is simulated for 14 weeks. We implement the following three algorithms: SingleAgent, MultiAgent, and MultiAgent+SurrogateRwd. Here the SingleAgent is the algorithm that trains a single agent that outputs all the three types of actions. The MultiAgent is the proposed MARL algorithm using the adherence as the learning reward signal for all three agents. The MultiAgent+SurrogateRwd is the proposed MARL algorithm using the surrogate reward functions. The full details of the algorithms are described in supplementary material.
We compare the proposed RL algorithms with a random policy, where $P(A_{w,d,t}^{\tiny\texttt{AYA}}=1)=P(A_{w,d}^{\tiny\texttt{CARE}}=1)=P(A_{w}^{\tiny\texttt{REL}}=1)\equiv 0.5$ in terms of the cumulative adherence improvement.
We also observe that all the algorithms can make more significant improvement over the random policy under a higher STE. SingleAgent takes longer to learn due to the larger number of parameters compared to MultiAgent. We also see an advantage of using surrogate rewards through an increased cumulative adherence at all levels of STE. Notice that for a low STE, the learning is slow, which is intuitive given that the the signal-to-noise ratio is low in such an environment. Additional ablation studies and analysis on the collaborating behavior is provided in the supplementary material.
@@ 6 Discussion
In this paper, we propose an MARL algorithm that effectively learns to optimize delivery of the ADAPTS-HCT digital interventions. While this presents a significant step towards preparing for the ADAPTS-HCT clinical trial, several challenges remain to be addressed. First, in the real-clinical trial, the participants are recruited incrementally with significant overlaps, whereas our dyadic environment assumes a simple sequential recruitment. Second, the clinical trial study emphasizes the need for after-study analysis, such as causal inference on treatment effects, which often requires smooth allocation functions [36]. Additionally, there is room to further improve algorithm performance. For example, our proposed algorithm pools data across dyads to reduce learning variance but does not account for heterogeneity across dyads. The algorithm may benefit from a more flexible pooling, e.g., a random effect model.
Appendix A Related Work

### v2

@@ Abstract
11footnotetext: Corresponding author: Ziping Xu, zipingxu@fas.harvard.edu
Medication adherence is critical for the recovery of adolescents and young adults (AYAs) who have undergone hematopoietic cell transplantation. However, maintaining adherence is challenging for AYAs after hospital discharge, who experience both individual (e.g. physical and emotional symptoms) and interpersonal barriers (e.g., relational difficulties with their care partner, who is often involved in medication management). To optimize the effectiveness of a three-component digital intervention targeting both members of the dyad as well as their relationship, we propose a novel Multi-Agent Reinforcement Learning (MARL) approach to personalize the delivery of interventions. By incorporating the domain knowledge, the MARL framework, where each agent is responsible for the delivery of one intervention component, allows for faster learning compared with a flattened agent. Evaluation using a dyadic simulator environment, based on real clinical data, shows a significant improvement in medication adherence (approximately 3%) compared to purely random intervention delivery. The effectiveness of this approach will be further evaluated in an upcoming trial.
Keywords:
Reinforcement Learning Dyadic Relationships Medication Adherence Digital Health
@@ 1 Introduction
The use of MARL decouples the decision processes of different intervention components, thus improving interpretability of the agent model design. This improved interpretability allows us to incorporate domain knowledge into the agent-specific algorithm designs to address challenge (2). To further accelerate learning, we propose a novel reward engineering method that learns a less noisy surrogate reward function for each component. Through evaluation in a carefully designed dyadic environment, we demonstrate both the superior performance of our proposed algorithm.
@@ 2 RL Framework and Domain Knowledge
HCT treatment is followed by an outpatient 100 days (rounded to 14 weeks) twice-daily medication regimen. Decision times within the 14 weeks are denoted by $(w,d,t)$ where $w\in\{1,\dots,14\}$ is the week index, $d\in\{1,\dots,7\}$ is the day index, and $t\in\{1,2\}$ is the decision window within a day.
@@
Action space. All actions are binary (deliver versus do not deliver intervention content); see Tables 1,2. When the current time $(d=1,t=1)$ is the first decision time on the first day of the week, the agent chooses three actions corresponding to all three interventions components. If the current time is the first time on a day after the first day of the week $(d>1,t=1)$, the agent chooses two actions corresponding to only the AYA intervention and the care partner components. At the second time on each day ($t=2$) the agent chooses one action corresponding to only the AYA intervention component.
Observation space. Apart from the dynamic action space, we collect observations about different components at different time scales (Table 2).
@@ 3 Domain Knowledge through Causal Diagram
Our algorithm design is guided by domain knowledge encoded as the causal diagram in Fig. 1. This diagram describes the scientific team’s understanding of the primary causal relationships between the variables in each component listed in Table 2.
These variables are selected by the scientific team that are believed to capture the major causal pathways from our interventions to the primary outcome. Specifically, the app burdens are included as it has been well established in the mobile health literature that the app burden is a key moderator of the digital intervention through mobile health apps [34]. Literature has shown that a supporting positive framing in a dyad is associated with better adherence [18, 7], which motivates the inclusion of the relationship quality. The dyadic relationship is highly associated with the psychological distress of the care partner. The later is the major challenge faced by the care partner as they operate the caregiving role [28].
We remark that the causal relationships are likely more complex and direct paths may exist between any two variables. However the scientific team believes that these other paths are likely to be less detectable given the noise in digital intervention data. We summarize the primary pathways that interventions can take to effect the AYA’s adherence in the following.
@@
    Game intervention. The game intervention $A_{w}^{\tiny\texttt{REL}}$ has two pathways by which it is expected to effect AYA’s adherence. First, $A_{w}^{\tiny\texttt{REL}}$ is expected to increase the AYA’s burden $B_{w,d,t}^{\tiny\texttt{AYA}}$ throughout the week $w$, which decreases the AYA’s adherence $R_{w,d,t}^{\tiny\texttt{AYA}}$ (blue arrows). Second, the game intervention $A_{w}^{\tiny\texttt{REL}}$ is expected to effect next week AYA’s adherence $R_{w+1,d,t}^{\tiny\texttt{AYA}}$ by improving the end of the week relationship quality $Y_{w}^{\tiny\texttt{REL}}$ (green arrows).
@@
    Care partner intervention. The care-partner intervention $A_{w,d}^{\tiny\texttt{CARE}}$ is expected to effect the AYA’s adherence indirectly. First, $A_{w,d}^{\tiny\texttt{CARE}}$ should decrease the care partner’s psychological distress $Y_{w,d}^{\tiny\texttt{CARE}}$. An decrease in psychological distress should increase the end of week relationship quality $Y_{w}^{\tiny\texttt{REL}}$ (yellow arrows). Second, $A_{w,d}^{\tiny\texttt{CARE}}$ should increase the care partner’s burden $B_{w,d}^{\tiny\texttt{CARE}}$, which should decrease the the end of week relationship quality $Y_{w}^{\tiny\texttt{REL}}$ (blue arrows).
@@ 4 Proposed Multi-Agent RL Approach
The conditional independence property observed from Fig. 1 motivates us to design a multi-agent RL (MARL) comprising three decoupled agents: the AYA agent, the care partner agent, and the relationship agent. Each agent makes decisions at different time scales for their own component.
@@ 4.1 Surrogate Reward Function Design Through Domain Knowledge
The effects of care partner intervention and the game intervention are highly delayed. The game intervention improves end of week relationship quality with a significant delayed effect onto the adherence in the next week. As shown in Fig. 1, the care partner intervention (positive messages for the care partner) is designed to mitigate the care partner’s psychological distress, which only has indirect and delayed effects on the AYA’s adherence. To address the this issue and accelerate learning, we engineer the reward function to account for the delayed effects. Similar reward engineering in the context of digital interventions is discussed in [35]. Our approach is distinct in that we explore the principles for incorporating domain knowledge to guide the reward design.
@@
Domain knowledge informed surrogate reward functions. We introduce the surrogate reward functions for the relationship agent and the care partner agent. As informed by Fig. 1, the delayed effect of the game intervention is through the relationship quality and the AYA burden. This motivates us to fit a linear model to predict the sum of medication adherence within week $w$, $\sum_{d=1}^{7}\sum_{t=1}^{2}R_{w,d,t}^{\tiny\texttt{AYA}}$, using $(1,Y_{w-1}^{\tiny\texttt{REL}},R_{w-1,7,2}^{\tiny\texttt{AYA}},{B}_{w,1,1}^{\tiny\texttt{AYA}},A_{w}^{\tiny\texttt{REL}},A_{w}^{\tiny\texttt{REL}}\cdot Y_{w-1}^{\tiny\texttt{REL}})$ as the covariates. To account for the delayed effect, we engineer the surrogate reward function for the relationship agent as: $r_{w}^{\tiny\texttt{REL}}=(1,Y_{w-1}^{\tiny\texttt{REL}},R_{w-1,7,2}^{\tiny\texttt{AYA}},{B}_{w,1,1}^{\tiny\texttt{AYA}},A_{w}^{\tiny\texttt{REL}},A_{w}^{\tiny\texttt{REL}}\cdot Y_{w-1}^{\tiny\texttt{REL}}){\bm{\beta}}^{\tiny\texttt{REL}}+\max_{a}(1,Y_{w}^{\tiny\texttt{REL}},R_{w,7,2}^{\tiny\texttt{AYA}},{B}_{w+1,1,1}^{\tiny\texttt{AYA}},a,a\cdot Y_{w}^{\tiny\texttt{REL}}){\bm{\beta}}^{\tiny\texttt{REL}},$ where ${\bm{\beta}}^{\tiny\texttt{REL}}\in\mathbb{R}^{5}$ are Bayesian linear regression estimates. The above reward yields a two-step greedy policy, which is a good enough approximation for the total sum of the medication adherence. We opt for a simple, linear model here because the bias trade-off is justified by the faster learning and reduction in noise.
@@
The design of the care partner agent is similar. A key observation is that the end of the week relationship quality blocks all the paths from the care partner variables to the next week’ AYA adherence. Additionally, we believe that the relationship quality only have positive effects. Thus, we fit a linear model to predict the end of week relationship quality $Y_{w+1}^{\tiny\texttt{REL}}$ using $(1,Y_{w,d}^{\tiny\texttt{CARE}},B_{w,d+1}^{\tiny\texttt{CARE}},Y_{w-1}^{\tiny\texttt{REL}},A_{w}^{\tiny\texttt{REL}},A_{w,d}^{\tiny\texttt{CARE}})$ as covariates. The surrogate reward function is: $r_{w,d}^{\tiny\texttt{CARE}}=(1,Y_{w,d}^{\tiny\texttt{CARE}},B_{w,d+1}^{\tiny\texttt{CARE}},Y_{w-1}^{\tiny\texttt{REL}},A_{w}^{\tiny\texttt{REL}},A_{w,d}^{\tiny\texttt{CARE}}){\bm{\beta}}^{\tiny\texttt{CARE}}$, where ${\bm{\beta}}^{\tiny\texttt{CARE}}\in\mathbb{R}^{5}$ are Bayesian linear regression estimates.
@@ 5 Results
Our environment is based on Roadmap 2.0 dataset [29] involving 171 dyads, each consisting of a patient undergone HCT (target person) and a care partner. Roadmap 2.0 provides daily positive psychology interventions to the care partner only. Roadmap 2.0 collects wearable devices data, for example, physical activity, and self-report data, for example, mood score.
We build upon previous work [13] that also uses the Roadmap 2.0 data, but primarily focuses on AYA and relationship intervention components. We extend the environment to include the care partner intervention component. Specifically, we fit a separate multi-variate linear model for each component’s outcome (ie., $R_{w,d,t}^{\tiny\texttt{AYA}},Y_{w,d}^{\tiny\texttt{CARE}},Y_{w}^{\tiny\texttt{REL}}$) in the dataset. These models simulate the user trajectories under no intervention. A complete description and code of the dyadic environment is provided in supplementary material  [^1] https://github.com/StatisticalReinforcementLearningLab/ADAPTS-HCT-AIME.
To simulate outcomes under treatments, we impute the treatment effects of the interventions and the effects of app burden, so the induced standard treatment effects (STE)  [^2] STE here is defined as the difference in the mean of the primary outcomes under the proposed intervention package and these under no intervention, which is further standardized by the standard deviation under no intervention. are around 0.15, 0.3, and 0.5. These STEs are commonly seen in behavioral science studies [3].
@@ 5.1 Cumulative Adherence Improvement
We simulate 25 dyads, the planned sample size in the upcoming pilot study, by sampling dyads sequentially with replacement from Roadmap 2.0 dataset. Each dyad is simulated for 14 weeks. We implement the following three algorithms: SingleAgent, MultiAgent, and MultiAgent+SurrogateRwd. Here the SingleAgent is the algorithm that trains a single agent that outputs all the three types of actions. The MultiAgent is the proposed MARL algorithm using the adherence as the learning reward signal for all three agents. The MultiAgent+SurrogateRwd is the proposed MARL algorithm using the surrogate reward functions.
We report the cumulative adherence improvement over the purely random policy in Figure 2. We observe that all the algorithms can make more significant improvement over the random policy under a higher STE. SingleAgent takes longer to learn due to the larger number of parameters compared to MultiAgent as it does not fully utilize the dyadic structure and the causal relationship between the components described in Figure 1. We also see an advantage of using surrogate rewards through an increased cumulative adherence at all levels of STE. Notice that for a low STE, the learning is slow, which is intuitive given that the the signal-to-noise ratio is low in such an environment. Additional ablation studies and analysis on the collaborating behavior is provided in the supplementary material.
@@ 6 Discussion
In this paper, we propose an MARL algorithm that effectively learns to optimize delivery of the ADAPTS-HCT digital interventions. While this presents a significant step towards preparing for the ADAPTS-HCT clinical trial, several challenges remain to be addressed. First, in the real-clinical trial, the participants are recruited incrementally with significant overlaps, whereas our dyadic environment assumes a simple sequential recruitment.
Acknowledgments. This research was supported by various grants: NIH/NIDA P50DA054039, NIH/NIBIB and OD P41EB028242, NIH/NIDCR UH3DE028723, and NIH/NIA 5P30AG073107-03 GY3 Pilots. Sung Won Choi is supported by NIH grants (K24HL156896 and R01CA249211 and UG1HL069330) and the Edith S Briskin/SKS Research Professorship and the Taubman medical Research Institute, which collectively support work related to GVHD.
Appendix 0.A Related Work